# Survey on Factuality in Large Language Models: Knowledge, Retrieval and Domain-Specificity

### v1

@@ 4.2.1 Model-level Causes
The model does not always retain knowledge from its training corpus. While it’s a misconception that LLMs “memorize” data, they do form representations of knowledge based on their training. However, they might not always recall specific, less-emphasized details from their training parameters, especially if such details were rare or not reinforced through multiple examples. For example, ChatGPT has pretrained with Wikipedia, but it still fail in answering some questions from NaturalQuestions (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017), which are constructed from Wikipedia (Wang et al., 2023a).
@@
While the model might possess relevant knowledge, it can sometimes fail to reason with it effectively to answer queries. Even if an LLM has the requisite knowledge to answer a question, it might fail to connect the dots or reason logically. For instance, ambiguity in the input (Liu et al., 2023a) can potentially lead to a failure in understanding by LLMs, consequently resulting in reasoning errors. In addition, Berglund et al. (2023) find the LLM traps in the reversal curse, for example, it knows that A is B’s mother but fail to answer who is B’s son. This is especially evident in complex multi-step reasoning tasks or when the model needs to infer based on a combination of facts (Tan et al., 2023b; Kotha et al., 2023).
4.2.2 Retrieval-level Causes
@@
LLMs can sometimes accept and propagate misinformation present in the retrieved data. This is especially concerning when the model encounters knowledge conflicts, where the retrieved information contradicts its pre-trained knowledge, or multiple retrieved documents contradict each other Li et al. (2015). For instance, Longpre et al. (2021) observed that the more irrelevant the evidence, the more likely the model is to rely on its intrinsic knowledge. Recent studies, such as Pan et al. (2023b), have also shown that LLMs are susceptible to misinformation attacks within the retrieval process.
@@
LLMs can be misled by irrelevant or distracting information in the retrieved data. For example, if the evidence mentions a ”Russian movie” and ”Directors”, the LLM might incorrectly infer that ”The director is Russian”. Luo et al. (2023a) highlighted this vulnerability, noting that LLMs can be significantly impacted by distracting retrieval results. They further proposed instruction tuning as a potential solution to enhance the model’s ability to sift through and leverage retrieval results more effectively.
@@
Even when the retrieved information is closely related to the query, LLMs can sometimes misunderstand or misinterpret it. While this might be less frequent when the retrieval process is optimized, it remains a potential source of error. For instance, in the ReAct study (Yao et al., 2023a), the rate of errors dropped significantly when the retrieval process was improved.
4.2.3 Inference-level Causes
@@
During the generation process, a minor error or deviation at the beginning can compound as the model continues to generate content. For instance, if an LLM misinterprets a prompt or starts with an inaccurate premise, the subsequent content can veer further from the truth (Zhang et al., 2023b; Varshney et al., 2023).
@@
LLMs are a product of their training data. If they’ve been exposed more frequently to certain types of content or phrasing, they might have a bias toward generating similar content, even when it’s not the most factual or relevant. This bias can be especially pronounced if the training data has imbalances or if certain factual scenarios are underrepresented (Felkner et al., 2023; Gallegos et al., 2023). The model’s outputs, in such cases, reflect its training exposure rather than objective factuality. For example, research by  Hossain et al. (2023) suggests that LLMs can correctly identify the gender of individuals who conform to the binary gender system, but they perform poorly when determining non-binary or neutral genders.
5 Enhancement
Figure 2: An overview of methods to enhance factuality in large language models. This encompasses three primary areas: enhancement techniques for pure LLMs, strategies for retrieval-augmented LLMs, and methods employed by domain-specific LLMs to boost their factual accuracy within their respective domains.
@@ 5 Enhancement
TABLE VII: Performance of Select Factuality Enhancement Methods. The table displays performance metrics on various datasets for both baseline models and their enhanced counterparts, denoted with a $\rightarrow$. Due to space constraints, only a subset of datasets, metrics, and models from each work is presented.
Reference Dataset Metrics Baselines $\rightarrow$ Theirs Dataset Metrics Baselines $\rightarrow$ Theirs
Li et al. (2022c) NQ EM 34.5 $\rightarrow$ 44.35 (T5 11B) GSM8K ACC 77.0 $\rightarrow$ 85.0 (ChatGPT)
Yu et al. (2023) NQ EM 20.9 $\rightarrow$ 28.0 (InstructGPT) TriviaQA EM 57.5 $\rightarrow$ 59.0 (InstructGPT)
WebQA EM 18.6 $\rightarrow$ 24.6 (InstructGPT)
Chuang et al. (2023) FACTOR News ACC 58.3 $\rightarrow$ 62.0 (LLaMa-7B) FACTOR News ACC 61.1 $\rightarrow$ 62.5 (LLaMa-13B)
FACTOR News ACC 63.8 $\rightarrow$ 65.4 (LLaMa-33B) FACTOR News ACC 63.6 $\rightarrow$ 66.2 (LLaMa-65B)
FACTOR Wiki ACC 58.6 $\rightarrow$ 62.2 (LLaMa-7B) FACTOR Wiki ACC 62.6 $\rightarrow$ 66.2 (LLaMa-13B)
FACTOR Wiki ACC 69.5 $\rightarrow$ 70.3 (LLaMa-33B) FACTOR Wiki ACC 72.2 $\rightarrow$ 72.4 (LLaMa-65B)
TruthfulQA %Truth * Info 32.4 $\rightarrow$ 44.6 (LLaMa-13B) TruthfulQA %Truth * Info 34.8 $\rightarrow$ 49.2 (LLaMa-65B)
Li et al. (2022b) TruthfulQA %Truth * Info 32.4 $\rightarrow$ 44.4 (LLaMa-13B) TruthfulQA %Truth * Info 31.7 $\rightarrow$ 36.7 (LLaMa-33B)
TruthfulQA %Truth * Info 34.8 $\rightarrow$ 43.4 (LLaMa-65B)
Li et al. (2023d) NQ ACC 46.6 $\rightarrow$ 51.3 (LLaMA-7B) TriviaQA ACC 89.6 $\rightarrow$ 91.1 (LLaMA-7B)
MMLU ACC 35.7 $\rightarrow$ 40.1 (LLaMA-7B) TruthfulQA %Truth * Info 32.5 $\rightarrow$ 65.1 (Alpaca)
TruthfulQA %Truth * Info 26.9 $\rightarrow$ 43.5 (LLaMa-7B) TruthfulQA %Truth * Info 51.5 $\rightarrow$ 74.0 (Vicuna)
Cohen et al. (2023b) LAMA F1 50.7 $\rightarrow$ 80.8 (ChatGPT) TriviaQA F1 56.2 $\rightarrow$ 82.6 (ChatGPT)
NQ F1 60.6 $\rightarrow$ 79.1 (ChatGPT) PopQA F1 65.2 $\rightarrow$ 85.4 (ChatGPT)
LAMA F1 42.5 $\rightarrow$ 79.3 (GPT-3) TriviaQA F1 46.7 $\rightarrow$ 77.2 (GPT-3)
NQ F1 52.0 $\rightarrow$ 78.0 (GPT-3) PopQA F1 43.7 $\rightarrow$ 77.4 (GPT-3)
Weller et al. (2023) TriviaQA QUIP 31.6 $\rightarrow$ 33.6 (ChatGPT) NQ QUIP 32.8 $\rightarrow$ 34.3 (ChatGPT)
HotpotQA QUIP 28.3 $\rightarrow$ 29.2 (ChatGPT) ELI5 QUIP 24.1 $\rightarrow$ 26.5 (ChatGPT)
TriviaQA EM 77.8 $\rightarrow$ 78.8 (ChatGPT) NQ EM 32.9 $\rightarrow$ 34.8 (ChatGPT)
HotpotQA F1 35.7 $\rightarrow$ 36.6 (ChatGPT) ELI5 R-L 22.7 $\rightarrow$ 21.7 (ChatGPT)
Dhuliawala et al. (2023) MultiSpanQA F1 39.0 $\rightarrow$ 48.0 (LLaMA 65B) - FactScore 55.9 $\rightarrow$ 71.4 (LLaMA 65B)
- Avg. # facts 16.6$\rightarrow$ 12.3 (LLaMA 65B)
Yao et al. (2023a) HotpotQA EM 28.7 $\rightarrow$ 35.1 (PaLM-540B) FEVER ACC 57.1 $\rightarrow$ 62.0 (PaLM-540B)
Jiang et al. (2023b) 2WikiMultihopQA EM 28.2 $\rightarrow$ 51 (ChatGPT) StrategyQA EM 72.9 $\rightarrow$ 77.3 (ChatGPT)
WikiAsp UniEval 47.1 $\rightarrow$ 53.4 (ChatGPT) ASQA EM 33.8 $\rightarrow$ 41.3 (ChatGPT)
ASQA-hint EM 40.1 $\rightarrow$ 46.2 (ChatGPT)
Izacard et al. (2022) MMLU ACC 42.4 $\rightarrow$ 56.3 (T5-770M) MMLU ACC 50.4 $\rightarrow$ 59.9 (T5-3B)
MMLU ACC 54 $\rightarrow$ 65.8 (T5-11B)
Shi et al. (2023) MMLU ACC 68.3 $\rightarrow$ 73.2 (Codex) NQ EM 40.6 $\rightarrow$ 45.5(Codex)
TriviaQA EM 73.6 $\rightarrow$ 77.3 (Codex)
@@
Table VII provides a summary of enhancement methods and their respective improvements over baseline LLMs. It’s essential to recognize that various research papers may employ distinct experimental settings, such as zero-shot, few-shot, or full settings. Consequently, when examining this table, it’s important to note that performance metrics for different methods, even when evaluating the same metric on the same dataset, may not be directly comparable.
5.1 On Standalone LLM Generation
@@
(2) Enhancing Factual Knowledge from Supervised Data (Sec 5.1.2): Examples in this category include supervised fine-tuning strategies (Chung et al., 2022; Zhou et al., 2023a) that focus on integrating structured knowledge from knowledge graphs (KGs) or making precise adjustments to model parameters (Li et al., 2023d).
@@
5.1.1 Pretraining-based
@@ 5.1.2 Supervised Finetuning
Yang et al. (2023b) presents a development framework for Knowledge Graph-enhanced Large Language Models (KGLLMs), drawing from established technologies. They detail various enhancement strategies, including a before-training enhancement that refines input quality by integrating factual data; a during-training enhancement that synergizes textual and structural knowledge, utilizing tools like graph neural networks and attention mechanisms; multi-task learning which focuses on knowledge-guided pre-training tasks to bolster the factual knowledge acquisition of LLMs; and a post-training enhancement that fine-tunes LLMs for domain-specific tasks using knowledge-rich data. Additionally, the significance of prompt learning in LLMs is highlighted, emphasizing the importance of choosing appropriate prompt templates. They also suggest knowledge graphs as a valuable resource for crafting these templates to harness domain-specific knowledge.
@@
Instead of directly finetuning the model, model edit is a more precise approach to enhance the model’s factuality. By editing specific areas that are related to the fact, the model can correctly express that fact without compromising other unrelated knowledge. Current editing methods (Yao et al., 2023b) can be categorized into weight-preserved and weight-modified paradigms. When modifying the model’s weight, KN (Dai et al., 2022) and ROME (Meng et al., 2022) first analyze representations to locate those underlying factual errors and then directly update the relevant weights. Meanwhile, KE (De Cao et al., 2021) and MEND (Mitchell et al., 2022a) employ a hypernetwork to learn the necessary weight changes. While effective, the robustness and generalization of directly updating weights remains an open question.
@@ 5.1.4 Novel Prompt
Dhuliawala et al. (2023) present Chain-of-Verification (CoVe) to reduce factual errors.The CoVe strategy involves the model initially crafting a response, subsequently formulating verification queries to assess its initial draft, independently responding to these queries to maintain unbiased answers, and ultimately producing a validated reply. The CoVe method encompasses four pivotal stages: 1) Drafting an initial reply based on a query using the LLM, 2) Generating verification questions from the query and initial answer to pinpoint potential errors, 3) Responding to each verification question and comparing these answers with the initial response to detect discrepancies, and 4) If discrepancies are found, producing a revised answer that integrates the verification outcomes. The entire procedure is executed by prompting the same LLM differently to achieve the intended results. Experiments show that Chain-of-Verification can reduce errors in diverse tasks, including list-based questions from Wikidata (Vrandečić and Krötzsch, 2014), closed book MultiSpanQA (Li et al., 2022a) and longform text generation (Min et al., 2023).
5.1.5 Decoding

### v3

@@ 4.2.1. Model-level Causes
The model does not always retain knowledge from its training corpus. While it’s a misconception that LLMs “memorize” data, they do form representations of knowledge based on their training. However, they might not always recall specific, less-emphasized details from their training parameters, especially if such details were rare or not reinforced through multiple examples. For example, ChatGPT has pretrained with Wikipedia, but it still fail in answering some questions from NaturalQuestions (Kwiatkowski et al., 2019) and TriviaQA (Joshi et al., 2017), which are constructed from Wikipedia (Wang et al., 2023b).
@@
While the model might possess relevant knowledge, it can sometimes fail to reason with it effectively to answer queries. Even if an LLM has the requisite knowledge to answer a question, it might fail to connect the dots or reason logically. For instance, ambiguity in the input (Liu et al., 2023d) can potentially lead to a failure in understanding by LLMs, consequently resulting in reasoning errors. In addition, Berglund et al. (2023) find the LLM traps in the reversal curse, for example, it knows that A is B’s mother but fail to answer who is B’s son. This is especially evident in complex multi-step reasoning tasks or when the model needs to infer based on a combination of facts (Tan et al., 2023b; Kotha et al., 2023).
4.2.2. Retrieval-level Causes
@@
LLMs can sometimes accept and propagate misinformation present in the retrieved data. This is especially concerning when the model encounters knowledge conflicts, where the retrieved information contradicts its pre-trained knowledge, or multiple retrieved documents contradict each other (Li et al., 2015). For instance, (Longpre et al., 2021) observed that the more irrelevant the evidence, the more likely the model is to rely on its intrinsic knowledge. Recent studies, such as (Pan et al., 2023b), have also shown that LLMs are susceptible to misinformation attacks within the retrieval process.
@@
LLMs can be misled by irrelevant or distracting information in the retrieved data. For example, if the evidence mentions a ”Russian movie” and ”Directors”, the LLM might incorrectly infer that ”The director is Russian”. (Luo et al., 2023a) highlighted this vulnerability, noting that LLMs can be significantly impacted by distracting retrieval results. They further proposed instruction tuning as a potential solution to enhance the model’s ability to sift through and leverage retrieval results more effectively.
@@
Even when the retrieved information is closely related to the query, LLMs can sometimes misunderstand or misinterpret it. While this might be less frequent when the retrieval process is optimized, it remains a potential source of error. For instance, in the ReAct study (Yao et al., 2023b), the rate of errors dropped significantly when the retrieval process was improved.
4.2.3. Inference-level Causes
@@
During the generation process, a minor error or deviation at the beginning can compound as the model continues to generate content. For instance, if an LLM misinterprets a prompt or starts with an inaccurate premise, the subsequent content can veer further from the truth (Zhang et al., 2023g; Varshney et al., 2023).
@@
LLMs are a product of their training data. If they’ve been exposed more frequently to certain types of content or phrasing, they might have a bias toward generating similar content, even when it’s not the most factual or relevant. This bias can be especially pronounced if the training data has imbalances or if certain factual scenarios are underrepresented (Felkner et al., 2023; Gallegos et al., 2023). The model’s outputs, in such cases, reflect its training exposure rather than objective factuality. For example, research by  (Hossain et al., 2023) suggests that LLMs can correctly identify the gender of individuals who conform to the binary gender system, but they perform poorly when determining non-binary or neutral genders.
5. Enhancement
Figure 2. An overview of methods to enhance factuality in large language models. This encompasses three primary areas: enhancement techniques for pure LLMs, strategies for retrieval-augmented LLMs, and methods employed by domain-specific LLMs to boost their factual accuracy within their respective domains.
@@ 5. Enhancement
Table 8. Performance of Selected Factuality Enhancement Methods. The table displays performance metrics on various datasets for both baseline models and their enhanced counterparts, denoted with a $\rightarrow$. Due to space constraints, only a subset of datasets, metrics, and models from each work is presented.
Reference Dataset Metrics Baselines $\rightarrow$ Theirs Dataset Metrics Baselines $\rightarrow$ Theirs Li et al. (2022a) NQ EM 34.5 $\rightarrow$ 44.35 (T5 11B) GSM8K ACC 77.0 $\rightarrow$ 85.0 (ChatGPT) GenRead (Yu et al., 2023) NQ EM 20.9 $\rightarrow$ 28.0 (InstructGPT) TriviaQA EM 57.5 $\rightarrow$ 59.0 (InstructGPT) WebQA EM 18.6 $\rightarrow$ 24.6 (InstructGPT) DoLa (Chuang et al., 2023) FACTOR News ACC 58.3 $\rightarrow$ 62.0 (LLaMa-7B) FACTOR News ACC 61.1 $\rightarrow$ 62.5 (LLaMa-13B) FACTOR News ACC 63.8 $\rightarrow$ 65.4 (LLaMa-33B) FACTOR News ACC 63.6 $\rightarrow$ 66.2 (LLaMa-65B) FACTOR Wiki ACC 58.6 $\rightarrow$ 62.2 (LLaMa-7B) FACTOR Wiki ACC 62.6 $\rightarrow$ 66.2 (LLaMa-13B) FACTOR Wiki ACC 69.5 $\rightarrow$ 70.3 (LLaMa-33B) FACTOR Wiki ACC 72.2 $\rightarrow$ 72.4 (LLaMa-65B) TruthfulQA %Truth * Info 32.4 $\rightarrow$ 44.6 (LLaMa-13B) TruthfulQA %Truth * Info 34.8 $\rightarrow$ 49.2 (LLaMa-65B) CD (Li et al., 2022b) TruthfulQA %Truth * Info 32.4 $\rightarrow$ 44.4 (LLaMa-13B) TruthfulQA %Truth * Info 31.7 $\rightarrow$ 36.7 (LLaMa-33B) TruthfulQA %Truth * Info 34.8 $\rightarrow$ 43.4 (LLaMa-65B) ITI (Li et al., 2023d) NQ ACC 46.6 $\rightarrow$ 51.3 (LLaMA-7B) TriviaQA ACC 89.6 $\rightarrow$ 91.1 (LLaMA-7B) MMLU ACC 35.7 $\rightarrow$ 40.1 (LLaMA-7B) TruthfulQA %Truth * Info 32.5 $\rightarrow$ 65.1 (Alpaca) TruthfulQA %Truth * Info 26.9 $\rightarrow$ 43.5 (LLaMa-7B) TruthfulQA %Truth * Info 51.5 $\rightarrow$ 74.0 (Vicuna) FactTune (Ovadia et al., 2023) Biographies ACC 74.8 $\rightarrow$ 83.1 (Llama-2-Chat) Medical QA ACC 63.6 $\rightarrow$ 68.2 (Llama-2-Chat) LM vs LM (Cohen et al., 2023b) LAMA F1 50.7 $\rightarrow$ 80.8 (ChatGPT) TriviaQA F1 56.2 $\rightarrow$ 82.6 (ChatGPT) NQ F1 60.6 $\rightarrow$ 79.1 (ChatGPT) PopQA F1 65.2 $\rightarrow$ 85.4 (ChatGPT) LAMA F1 42.5 $\rightarrow$ 79.3 (GPT-3) TriviaQA F1 46.7 $\rightarrow$ 77.2 (GPT-3) NQ F1 52.0 $\rightarrow$ 78.0 (GPT-3) PopQA F1 43.7 $\rightarrow$ 77.4 (GPT-3) Weller et al. (2023) TriviaQA QUIP 31.6 $\rightarrow$ 33.6 (ChatGPT) NQ QUIP 32.8 $\rightarrow$ 34.3 (ChatGPT) HotpotQA QUIP 28.3 $\rightarrow$ 29.2 (ChatGPT) ELI5 QUIP 24.1 $\rightarrow$ 26.5 (ChatGPT) TriviaQA EM 77.8 $\rightarrow$ 78.8 (ChatGPT) NQ EM 32.9 $\rightarrow$ 34.8 (ChatGPT) HotpotQA F1 35.7 $\rightarrow$ 36.6 (ChatGPT) ELI5 R-L 22.7 $\rightarrow$ 21.7 (ChatGPT) Chain-of-Verification (Dhuliawala et al., 2023) MultiSpanQA F1 39.0 $\rightarrow$ 48.0 (LLaMA 65B) - FactScore 55.9 $\rightarrow$ 71.4 (LLaMA 65B) - Avg. # facts 16.6$\rightarrow$ 12.3 (LLaMA 65B) ReAct (Yao et al., 2023b) HotpotQA EM 28.7 $\rightarrow$ 35.1 (PaLM-540B) FEVER ACC 57.1 $\rightarrow$ 62.0 (PaLM-540B) Jiang et al. (2023a) 2WikiMultihopQA EM 28.2 $\rightarrow$ 51 (ChatGPT) StrategyQA EM 72.9 $\rightarrow$ 77.3 (ChatGPT) WikiAsp UniEval 47.1 $\rightarrow$ 53.4 (ChatGPT) ASQA EM 33.8 $\rightarrow$ 41.3 (ChatGPT) ASQA-hint EM 40.1 $\rightarrow$ 46.2 (ChatGPT) Atlas (Izacard et al., 2022) MMLU ACC 42.4 $\rightarrow$ 56.3 (T5-770M) MMLU ACC 50.4 $\rightarrow$ 59.9 (T5-3B) MMLU ACC 54 $\rightarrow$ 65.8 (T5-11B) REPLUG (Shi et al., 2023b) MMLU ACC 68.3 $\rightarrow$ 73.2 (Codex) NQ EM 40.6 $\rightarrow$ 45.5(Codex) TriviaQA EM 73.6 $\rightarrow$ 77.3 (Codex) Self-RAG (Asai et al., 2023) TriviaQA ACC 30.0 $\rightarrow$ 72.4 (LLaMA2-7B) ALCE-ASQA EM 15.2 $\rightarrow$ 30.0 (LLaMA2-13B) TriviaQA ACC 30.2 $\rightarrow$ 74.5 (LLaMA2-7B) ALCE-ASQA EM 16.3 $\rightarrow$ 31.7 (LLaMA2-13B)
@@
Table 8 provides a summary of enhancement methods and their respective improvements over baseline LLMs. It’s essential to recognize that various research papers may employ distinct experimental settings, such as zero-shot, few-shot, or full settings. Consequently, when examining this table, it’s important to note that performance metrics for different methods, even when evaluating the same metric on the same dataset, may not be directly comparable.
5.1. On Standalone LLM Generation
@@
(2) Enhancing Factual Knowledge from Supervised Data (Sec 5.1.2): Examples in this category include supervised fine-tuning strategies (Chung et al., 2022; Zhou et al., 2023a) focus on finetuning with labelled data or integrating structured knowledge such as knowledge graphs (KGs) or making precise adjustments to model parameters (Li et al., 2023d).
@@
Some work (Liu et al., 2023c; Sun et al., 2023b) aim for improve the factuality of large multi-modal models, we choose to not emphasize them.
5.1.1. Pretraining-based
@@ 5.1.2. Supervised Finetuning
Yang et al. (2023a) presents a development framework for Knowledge Graph-enhanced Large Language Models (KGLLMs), drawing from established technologies. They detail various enhancement strategies, including a before-training enhancement that refines input quality by integrating factual data; a during-training enhancement that synergizes textual and structural knowledge, utilizing tools like graph neural networks and attention mechanisms; multi-task learning which focuses on knowledge-guided pre-training tasks to bolster the factual knowledge acquisition of LLMs; and a post-training enhancement that fine-tunes LLMs for domain-specific tasks using knowledge-rich data. Additionally, the significance of prompt learning in LLMs is highlighted, emphasizing the importance of choosing appropriate prompt templates. They also suggest knowledge graphs as a valuable resource for crafting these templates to harness domain-specific knowledge.
FactTune (Ovadia et al., 2023) aims to use Reinforcement Learning to improve the factuality of LLM, so scoring the factuality of the content generated by the model is needed. However, human labels are too costly, so authors use two methods to check the factuality of the content generated by LLM, namely 1) Reference-based, which uses external knowledge base retrieval and Factscore (Min et al., 2023) method to score; 2) Reference-free, which uses Another LLM’s confidence to score. Then, they utilized the Directly Preference Optimization (Rafailov et al., 2023) algorithm to optimize the model. 50% and 20+error rate reductions were observed in Llama’s bio generation tasks and medical QA respectively.
@@
Instead of directly finetuning the model, model edit is a more precise approach to enhance the model’s factuality. By editing specific areas that are related to the fact, the model can correctly express that fact without compromising other unrelated knowledge. Current editing methods (Yao et al., 2023a) can be categorized into weight-preserved and weight-modified paradigms. When modifying the model’s weight, KN (Dai et al., 2022) and ROME (Meng et al., 2022) first analyze representations to locate those underlying factual errors and then directly update the relevant weights. Meanwhile, KE (De Cao et al., 2021) and MEND (Mitchell et al., 2022a) employ a hypernetwork to learn the necessary weight changes. While effective, the robustness and generalization of directly updating weights remains an open question.
@@ 5.1.4. Novel Prompt
Dhuliawala et al. (2023) present Chain-of-Verification (CoVe) to reduce factual errors.The CoVe strategy involves the model initially crafting a response, subsequently formulating verification queries to assess its initial draft, independently responding to these queries to maintain unbiased answers, and ultimately producing a validated reply. The CoVe method encompasses four pivotal stages: 1) Drafting an initial reply based on a query using the LLM, 2) Generating verification questions from the query and initial answer to pinpoint potential errors, 3) Responding to each verification question and comparing these answers with the initial response to detect discrepancies, and 4) If discrepancies are found, producing a revised answer that integrates the verification outcomes. The entire procedure is executed by prompting the same LLM differently to achieve the intended results. Experiments show that Chain-of-Verification can reduce errors in diverse tasks, including list-based questions from Wikidata (Vrandečić and Krötzsch, 2014), closed book MultiSpanQA (Li et al., 2022c) and longform text generation (Min et al., 2023).
5.1.5. Decoding